# WhiSPA: Semantically and Psychologically Aligned Whisper with Self-Supervised Contrastive and Student-Teacher Learning

### v2

@@ A.3 Annotations
The documentation accompanying the iHiTOP interview dataset was utilized to report the coverage of its domains, demographic information, and other relevant details. The dataset’s focus on structured psychological interviews and its linguistic properties were described in the paper to contextualize its relevance to this research. This information was presented to ensure transparency and reproducibility. The WTC dataset assessed PTSD symptom severity and related constructs, including anxiety and depression, using English-language data from WTC emergency responders. Linguistic features such as RoBERTa-large embeddings, n-grams, and LDA topics were used to analyze behavioural patterns alongside closed-vocabulary features like pronouns and death-related terms (LIWC-22). The development dataset included 1,437 participants, and the prospective dataset included 346, with a mean age of 58 years, predominantly male (93% and 91%, respectively) and white (54% and 49%). The analysis emphasized language markers of stress, anxiety, and trauma while reflecting on participants’ experiences of 9/11. Ethical safeguards, including IRB approval, informed consent, and automated anonymization, ensured compliance. While comprehensive in its linguistic and demographic scope, the study was limited to English speakers and WTC responders, constraining generalizability.
@@ A.4 Quantitative Analysis
PsychEmb’s lower correlations in Figure 8 should not be mistaken for poor performance. With only 10 dimensions, PsychEmb representations achieve a staggering $24$ and $22$ Pearson points on INT and DIS respectively, emphasizing its validity as the psychological teacher. WhiSPA’s consistent improvement over the audio models is attributed to the semantic and psychological dimensions that SBERT and PsychEmb offer. Notably, WhiSPA exemplifies drastic improvements in prediction accuracy for VAL, INT, THT, and PCL compared to Whisper-384.
@@
While WhiSPA demonstrates substantial advancements, surpassing even its text-based LM teacher, SBERT-384, it remains inherently constrained by the representational capacity of the teacher model. If the teacher’s capabilities are limited, these deficiencies inevitably carry over to the student, even after distillation. This is evident in the ARO column, where arousal —- an affective dimension -— is more accurately conveyed through acoustic cues. However, WhiSPA struggles to capture and preserve the acoustic information, instead predominantly aligning with the semantic representations provided by SBERT, thus limiting its ability to fully represent the nuanced affective content inherent in speech.
@@
Beyond demonstrating superior alignment with established PTSD markers, Table 6 highlights WhiSPA’s enhanced sensitivity to psychologically meaningful language patterns. 6(a) shows that n-grams reflecting personal experiences, self-referential content (e.g., first-person pronouns), and negative affective states correlate more strongly with WhiSPA’s predictions than with those of Whisper. WhiSPA appears better attuned to indicators of psychological distress, anxiety, and trauma symptoms—-an advantage likely stemming from the contrastive alignment objective with text-based representations. The model’s capacity to detect nuanced emotional and cognitive expressions in spoken language is further supported by its higher effect sizes on known PTSD-relevant n-grams, underscoring that semantically oriented embeddings can bolster the recognition of clinically significant markers in audio data.
@@
Meanwhile, 6(b) points to a distinctive negative association between WhiSPA’s predicted severity scores and n-grams referencing positive affect or social relationships. This result suggests that the same semantically focused latent space that amplifies negative or distress-related terms also filters out language tied to more adaptive or supportive experiences. In practical terms, such an effect could be advantageous for screening or early detection: positive affect or relational talk might serve as a buffer or resilience indicator, thereby inversely correlating with predicted symptom severity. Taken together, these findings highlight the unique strength of WhiSPAs in capturing a wide spectrum of psychologically relevant linguistic markers, surpassing the granularity offered by audio models alone.
Figure 8: WhiSPA Closes the Semantic/Psychological Representation Gap. WhiSPA consistently outperforms every baseline audio model and, in some cases, even exceeds the performance of the text-based language model teacher.

### v4

@@ A.3 Annotations
The documentation accompanying the iHiTOP interview dataset was utilized to report the coverage of its domains, demographic information, and other relevant details. The dataset’s focus on structured psychological interviews and its linguistic properties were described in the paper to contextualize its relevance to this research. This information was presented to ensure transparency and reproducibility. The WTC dataset assessed PTSD symptom severity and related constructs, including anxiety and depression, using English-language data from WTC emergency responders in the Stony Brook Health Program. The WTC dataset assessed PTSD symptom severity and related constructs, including anxiety and depression, using English-language data from WTC emergency responders. The development dataset included 1,437 participants, and the prospective dataset included 346, with a mean age of 58 years, predominantly male (93% and 91%, respectively) and white (54% and 49%). The analysis emphasized language markers of stress, anxiety, and trauma while reflecting on participants’ experiences of 9/11. Ethical safeguards, including IRB approval, informed consent, and automated anonymization, ensured compliance. While comprehensive in its linguistic and demographic scope, the study was limited to English speakers and WTC responders, constraining generalizability.
@@ A.4 Quantitative Analysis
PsychEmb’s lower correlations in Figure 4 should not be mistaken for poor performance. With only 10 dimensions, PsychEmb representations achieve a staggering $24$ and $22$ Pearson points on INT and DIS respectively, emphasizing its validity as the psychological teacher. WhiSPA’s consistent improvement over the audio models is attributed to the semantic and psychological dimensions that SBERT and PsychEmb offer. Notably, WhiSPA exemplifies drastic improvements in prediction accuracy for all outcomes compared to Whisper.
@@
While WhiSPA demonstrates substantial advancements, surpassing even its text-based LM teacher, SBERT-1024, it remains inherently constrained by the representational capacity of the teacher model. If the teacher’s capabilities are limited, these deficiencies inevitably carry over to the student, even after distillation. This is evident in the ARO column, where arousal —- an affective dimension -— is more accurately conveyed through acoustic cues. However, WhiSPA struggles to capture and preserve the acoustic information, instead predominantly aligning with the semantic representations provided by SBERT, thus limiting its ability to fully represent the nuanced affective content inherent in speech.
@@
Beyond demonstrating superior alignment with established PTSD markers, Table 7 highlights WhiSPA’s enhanced sensitivity to psychologically meaningful language patterns. 7(a) shows that n-grams reflecting personal experiences, self-referential content (e.g., first-person pronouns), and negative affective states correlate more strongly with WhiSPA’s predictions than with those of Whisper. WhiSPA appears better attuned to indicators of psychological distress, anxiety, and trauma symptoms—-an advantage likely stemming from the contrastive alignment objective with text-based representations. The model’s capacity to detect nuanced emotional and cognitive expressions in spoken language is further supported by its higher effect sizes on known PTSD-relevant n-grams, underscoring that semantically oriented embeddings can bolster the recognition of clinically significant markers in audio data.
@@
Meanwhile, 7(b) points to a distinctive negative association between WhiSPA’s predicted severity scores and n-grams referencing positive affect or social relationships. This result suggests that the same semantically focused latent space that amplifies negative or distress-related terms also filters out language tied to more adaptive or supportive experiences. In practical terms, such an effect could be advantageous for screening or early detection: positive affect or relational talk might serve as a buffer or resilience indicator, thereby inversely correlating with predicted symptom severity. Taken together, these findings highlight the unique strength of WhiSPAs in capturing a wide spectrum of psychologically relevant linguistic markers, surpassing the granularity offered by audio models alone.